OpenStudy (anonymous):

simplify each radical expression. Use absolute value when necessary. 1) square root of a^4b^6 2) cube root of 216y^8

OpenStudy (anonymous):

\[\sqrt{a ^{4}b ^{6}}\] \[\sqrt[3]{216y ^{8}}\]

OpenStudy (anonymous):

@Callisto can you help?

OpenStudy (callisto):

\[\sqrt{a^4b^6}= (a^4b^6)^{\frac{1}{2}}=...\]Cube root is just 1/3 instead of 1/2

OpenStudy (anonymous):

so a^2b^3? and 6y^2?

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

thank you!!

OpenStudy (callisto):

Hmm.. Wait... \[\sqrt{a^2}=|a|\]

OpenStudy (anonymous):

wait what would that be then? \[\left| a^2 \right|\left| b^3 \right|\]?

OpenStudy (callisto):

That's just \[|a^2b^3|\]

OpenStudy (anonymous):

oh ok and then the other one would be \[6\left| y^2 \right|\]?

OpenStudy (callisto):

Nope.. That's cube root.. not the same as square root!!

OpenStudy (anonymous):

oh so it would just be 6y^2?

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

alright i think im done for the night thank you so much!!!

OpenStudy (callisto):

You're welcome :|

OpenStudy (anonymous):

i actually understand this stuff better now :D

OpenStudy (callisto):

Hmm.. I still have doubt about the first question here... :|

OpenStudy (anonymous):

its ok i can get help tomorrow then tell you so you'll know

OpenStudy (callisto):

Okay! Thanks :D Good night!!

OpenStudy (anonymous):

good night thanks for everything!

OpenStudy (callisto):

Wiki says it's \(\sqrt{x^2}=|x|\)~ Welcome :)